OpenStudy (anonymous):

Solve: -6 is less then or equal to 2x+1/4 is less then or equal to 6.

OpenStudy (lynfran):

\[0\le x \le \frac{ -1 }{ 8 }\]

OpenStudy (lynfran):

\[-6\le 2x+\frac{ 1 }{ 4 }\le 6\]\[0\le6+2x+\frac{ 1 }{ 4 }-6\le0\]\[0\le2x+\frac{ 1 }{ 4 }\le0\]\[0lex \le \frac{ -1 }{ 4 }\]\[0\le x \le \frac{ -1 }{ 8 }\]

OpenStudy (lynfran):

the second to last shd be \[0\le 2x \le \frac{ -1 }{ 4 }\]